# Self-eXplainable AI for Medical Image Analysis: A Survey and New Outlooks

### v1

@@ Abstract
The increasing demand for transparent and reliable models, particularly in high-stakes decision-making areas such as medical image analysis, has led to the emergence of eXplainable Artificial Intelligence (XAI). Post-hoc XAI techniques, which aim to explain black-box models after training, have been controversial in recent works concerning their fidelity to the models’ predictions. In contrast, Self-eXplainable AI (S-XAI) offers a compelling alternative by incorporating explainability directly into the training process of deep learning models. This approach allows models to generate inherent explanations that are closely aligned with their internal decision-making processes. Such enhanced transparency significantly supports the trustworthiness, robustness, and accountability of AI systems in real-world medical applications. To facilitate the development of S-XAI methods for medical image analysis, this survey presents an comprehensive review across various image modalities and clinical applications. It covers more than 200 papers from three key perspectives: 1) input explainability through the integration of explainable feature engineering and knowledge graph, 2) model explainability via attention-based learning, concept-based learning, and prototype-based learning, and 3) output explainability by providing counterfactual explanation and textual explanation. Additionally, this paper outlines the desired characteristics of explainability and existing evaluation methods for assessing explanation quality. Finally, it discusses the major challenges and future research directions in developing S-XAI for medical image analysis.
@@ 1 Introduction
This paper presents the first systematic review of S-XAI for medical image analysis, covering methodology, medical applications, and evaluation metrics, while also offering an in-depth discussion on challenges and future directions. Although there is a wealth of literature on medical XAI surveys [17, 2, 18, 19, 20, 21, 22] that deliver valuable insights, none have focused specifically on a comprehensive review of S-XAI methods applied to medical image analysis. We analyze more than 200 papers published from 2018 to September 2024, sourced from the proceedings of NeurIPS, ICLR, ICML, AAAI, CVPR, ICCV, ECCV, and MICCAI as well as top-tier journals in the field, including Nature Medicine, Transactions on Pattern Analysis and Machine Intelligence, Transactions on Medical Imaging, Medical Image Analysis, or those cited in related works. The statistics of research articles using keywords self-explainable, medical, image on Google Scholar are presented in Fig. 2, which reveal two key observations: 1) there has been a significant and consistent increase in research papers focused on self-explainable AI over the years, indicating growing interest and emphasis within the research community; 2) nearly half of the total research papers (green bars) are dedicated to applying S-XAI techniques in medical imaging (blue bars), highlighting the vital importance of S-XAI in the medical field.
@@
    Novel Scope of XAI Survey: As an emerging XAI method that actively offers explainability from the model itself, S-XAI is attracting growing attention from the research community. This work represents the first comprehensive survey on this topic.
@@
    Systematic Review of Methods: We present a novel taxonomy of relevant papers and review them based on input explainability, model explainability, and output explainability. This offers insights into potential technical innovations for S-XAI methods.
@@
    Thorough Overview of Applications: We overview various applications across different anatomical locations and modalities in current S-XAI research. This illustrates the ongoing development of S-XAI technologies in medical image practices, serving as a reference for future applications in diverse scenarios.
@@
    In-depth Discussion of Challenges and Future Work: We discuss the key challenges and look forward to the promising future directions. This highlights current shortcomings and identifies new opportunities for researchers to drive further advancements.
@@ 3.1 Explainable Feature Engineering
Feature engineering focuses on transforming raw images into a more useful set of human-interpretable features. This process is crucial for traditional machine learning methods to achieve accurate predictions, but it can be time-consuming and demands significant domain expertise. In contrast, deep learning models automatically extract features from raw images, simplifying the manual crafting process but often resulting in reduced interpretability. A promising approach to enhance input explainability is to incorporate explainable feature engineering into deep learning, which injects domain knowledge into the model, as shown in Fig. 4(a). This integration enhances the model’s interpretability by ensuring that the learned features are relevant and meaningful for clinical applications. Ultimately, this method improves model performance and offers valuable insights into the decision-making process.
A common strategy in explainable feature engineering is to combine both handcrafted and deep features from an input image to make final predictions [24, 25]. For example, Kapse et al.[24] introduce a self-interpretable multiple instance learning (SI-MIL) framework that simultaneously learns from deep image features and handcrafted morphometric and spatial descriptors. They assess the local and global interpretability of SI-MIL through statistical analysis, a user study, and key interpretability criteria. Another line of approach involves incorporating interpretable clinical variables as additional inputs alongside the images, often utilizing multimodal learning techniques [26, 27]. For instance, Xiang et al.[26] introduce OvcaFinder, an interpretable model that combines deep learning predictions from ultrasound images with Ovarian–Adnexal Reporting and Data System scores provided by radiologists, as well as routine clinical variables for diagnosing ovarian cancer. This approach enhances diagnostic accuracy and explains the impact of key features and regions on the prediction outcomes.
@@ 3.2 Knowledge Graph
Figure 4: Input explainability that incorporates (a) explainable feature engineering (b) knowledge graph as additional inputs.
A knowledge graph (KG) is a structured representation of factual knowledge that captures relationships between entities in a specific area. It provides a way to organize and represent knowledge in a semantically rich and interconnected manner and plays a crucial role in enhancing the interpretability of S-XAI models. Recently, integrating structured domain knowledge into downstream tasks has attracted significant attention of both industry and academia[28, 29, 30]. This growing interest stems from the recognition that leveraging domain knowledge can greatly improve the performance and effectiveness of various applications. As shown in Fig 4(b), regarding medical imaging analysis, the utilization of KG can be broadly categorized into three categories: 1) prior KG, which serves as a foundational resource that gathers existing domain expertise and established medical knowledge; 2) data KG, which is derived from the analysis of large-scale medical imaging datasets; and 3) hybrid KG, which combines the strengths of both prior and data KGs for medical image analysis.

### v2

@@ Abstract
The increasing demand for transparent and reliable models, particularly in high-stakes decision-making areas such as medical image analysis, has led to the emergence of eXplainable Artificial Intelligence (XAI). Post-hoc XAI techniques, which aim to explain black-box models after training, have raised concerns about their fidelity to model predictions. In contrast, Self-eXplainable AI (S-XAI) offers a compelling alternative by incorporating explainability directly into the training process of deep learning models. This approach allows models to generate inherent explanations that are closely aligned with their internal decision-making processes, enhancing transparency and supporting the trustworthiness, robustness, and accountability of AI systems in real-world medical applications. To facilitate the development of S-XAI methods for medical image analysis, this survey presents a comprehensive review across various image modalities and clinical applications. It covers more than 200 papers from three key perspectives: 1) input explainability through the integration of explainable feature engineering and knowledge graph, 2) model explainability via attention-based learning, concept-based learning, and prototype-based learning, and 3) output explainability by providing textual and counterfactual explanations. This paper also outlines desired characteristics of explainability and evaluation methods for assessing explanation quality, while discussing major challenges and future research directions in developing S-XAI for medical image analysis.
@@ 1 Introduction
This paper presents the first systematic review of S-XAI for medical image analysis, covering methodology, medical applications, and evaluation metrics, while also offering an in-depth discussion on challenges and future directions. Although there is a wealth of literature on medical XAI surveys [17, 2, 18, 19, 20, 21, 22] that deliver valuable insights, none have focused specifically on a comprehensive review of S-XAI methods applied to medical image analysis. We analyze more than 200 papers published from 2018 to October 2024, sourced from the proceedings of NeurIPS, ICLR, ICML, AAAI, CVPR, ICCV, ECCV, and MICCAI as well as top-tier journals in the field, including Nature Medicine, Transactions on Pattern Analysis and Machine Intelligence, Transactions on Medical Imaging, Medical Image Analysis, or those cited in related works. The statistics of research articles using keywords self-explainable, medical, image on Google Scholar are presented in Fig. 2, which reveal two key observations: 1) there has been a significant and consistent increase in research papers focused on self-explainable AI over the years, indicating growing interest and emphasis within the research community; 2) nearly half of the total research papers (green bars) are dedicated to applying S-XAI techniques in medical imaging (blue bars), highlighting the vital importance of S-XAI in the medical field.
@@
    Novel scope of XAI survey: As an emerging XAI method that actively offers explainability from the model itself, S-XAI is attracting growing attention from the research community. This work serves as the first comprehensive survey on this topic.
@@
    Systematic review of methods: We introduce a novel taxonomy of relevant papers and review them based on input explainability, model explainability, and output explainability. This offers insights into potential technical innovations for S-XAI methods.
@@
    Thorough overview of applications: We overview various downstream applications across different anatomical locations and modalities in current S-XAI research. This illustrates the ongoing development of S-XAI technologies in medical image analysis, serving as a reference for future applications in diverse contexts.
@@
    In-depth discussion of challenges and future work: We discuss the key challenges and look forward to the promising future directions. This underscores existing shortcomings and identifies new outlooks for researchers to drive further advancements.
@@ 3.1 Explainable Feature Engineering
Feature engineering aims to transform raw images into a more useful set of human-interpretable features [24], which is essential for traditional machine learning. However, deep learning models automatically extract features from raw images, simplifying the handcraft process but often sacrificing interpretability. To enhance input explainability, it is beneficial to combine explainable feature engineering with deep learning, as shown in Fig. 4A. This integration enhances models’ interpretability by injecting domain knowledge into models and ensuring that the learned features are relevant and meaningful for clinical applications.
A common strategy is to combine both deep and handcrafted features from an input image to make final predictions [25, 26]. For example, Kapse et al.[25] introduce a self-interpretable multiple instance learning framework that simultaneously learns from deep image features and handcrafted morphometric and spatial descriptors. Local and global interpretability are demonstrated through quantitative and qualitative benchmarking. Another line of approach involves incorporating interpretable clinical variables as additional inputs alongside images, often utilizing multimodal learning techniques [27, 28]. For instance, Xiang et al.[27] introduce OvcaFinder, an interpretable model that combines deep learning predictions from ultrasound images with Ovarian–Adnexal Reporting and Data System scores provided by radiologists, as well as routine clinical variables for diagnosing ovarian cancer. This approach enhances diagnostic accuracy and explains the impact of key features and regions on the prediction outcomes.
@@ 3.2 Knowledge Graph
Figure 4: Input explainability that incorporates explainable feature engineering and knowledge graphs as additional inputs.
A knowledge graph (KG) is a structured representation of factual knowledge that captures relationships between entities in the real world [29]. Recently, there has been increasing interest in integrating structured domain knowledge into downstream applications [30, 31, 32], with the recognition that leveraging domain knowledge can greatly improve the performance and self-explainability of deep learning models. Regarding medical image analysis, the types of KG can be broadly categorized into three categories: 1) prior KG, which serves as a foundational resource that gathers existing domain expertise; 2) data KG, which is derived from the analysis of large-scale medical image datasets; and 3) hybrid KG, which combines prior KG and data KG, as illustrated in Fig. 4B.